OpenStudy (anonymous):

Find all real solutions of the equation(if there is a no solutions, enter NO SOLUTION) root3(16a-40)=2

OpenStudy (anonymous):

\[\sqrt[3]{16a - 40}=2\] ?

OpenStudy (anonymous):

Or is it \[\large{\sqrt{3(16a -40) }= 2}\]

OpenStudy (anonymous):

Tell me Ep471_08